# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Integrating GenAI into Collaborative Newsroom Routines

### v2

@@ Abstract.
Generative AI (GenAI) is reshaping work, but adoption remains largely individual and experimental rather than integrated into collaborative routines. Whether GenAI can move from individual use to collaborative work is a critical question for future organizations. Journalism offers a compelling site to examine this shift: individual journalists have already been disrupted by GenAI tools; yet newswork is inherently collaborative relying on shared routines and coordinated workflows. We conducted 27 interviews with newsrooms managers, editors, and front-line journalists in China. We found that journalists frequently used GenAI to support daily tasks, but value alignment was safeguarded mainly through individual discretion. At the organizational level, GenAI use remained disconnected from team workflows, hindered by structural barriers and cultural reluctance to share practices. These findings underscore the gap between individual and collective adoption, pointing to the need for accounting for organizational structures, cultural norms, and workflow integration when designing GenAI for collaborative work.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Organizations today face a profound question: How should today’s organizations adopt generative artificial intelligence (GenAI), and how could GenAI become a routine part of collaborative work in future organizations? In 2025, Stanford HAI and Google DeepMind (Stanford Institute for Human-Centered Artificial Intelligence and Google DeepMind, 2025) invited researchers worldwide to explore future collaborative work within organizations, emphasizing that “AI is transforming how we work, make decisions, and collaborate. But while our tools are evolving fast, our organizational models are overdue for reinvention.” Despite its transformative potential, GenAI has entered most workplaces slowly and cautiously. Early adoption often takes the form of individual experiments and tentative trials, rather than sweeping integration led by organizations (Xiao et al., 2025b; Wu, 2024). Across industries, organizations are grappling with a familiar tension in technology adoption: the promise of innovation collides with the uncertainty of organizational change, as leaders and practitioners question how GenAI will reshape collaborative routines and align, or misalign, with teamwork norms (Gartner, 2024). In this study, we investigate how GenAI might move beyond fragmented individual use to become an integral part of organizational collaboration, and what challenges and opportunities this transition creates for reconfiguring collaborative work around GenAI in organizations.
@@
In this paper, we use the newsroom as our primary empirical site to examine what collaborative work with GenAI entails in practice. News production is rarely the effort of a single individual: front-line journalists gather information, draft stories, and exchange leads with peers; their work is then reviewed, edited, and sometimes rewritten by editors; and senior managers provide strategic oversight, shape policies, and liaise with external stakeholders (Duffy, 2021; Xiao et al., 2025b; Tuchman, 1973). Collaboration in newsrooms therefore spans both inter-role and intra-role interactions, anchored in shared routines and collective values such as editorial responsibility and mutual trust (Tuchman, 1973; Carlson, 2009; Robinson, 2011). We distinguish between three key roles around collaborative newswork in our study: front-line journalists, who are primarily responsible for story development and on-the-ground reporting; editors, who oversee content refinement, ensure editorial consistency, and coordinate team workflows; and senior managers, the newsroom leaders responsible for organizational decision-making and long-term strategy. While journalism scholarship often uses the umbrella term journalists to refer broadly to various news practitioners (Duffy, 2021), we adopt this convention as well. In our analysis of how GenAI reshapes collaborative practices and team dynamics, we therefore attend to collaboration within roles as well as across roles.
@@
While these roles structure collaboration in conventional newsroom routines, the rise of GenAI introduces new uncertainties about how such collaboration should unfold. Recently, on the one hand, many individual journalists are experimenting with GenAI in tentative, ad hoc ways (Liang et al., 2025), often relying on personal judgment to ensure that its use remains consistent with professional norms (Wu, 2024). On the other hand, major news organizations have started to articulate institutional responses to the rise of GenAI, for example, China Media Group (CMG) is developing proprietary GenAI systems for its news production (China Press, Publication, Radio, Film and Television Journal, 2024; Lang, 2024), while The British Broadcasting Corporation (BBC) has issued formal usage guidelines for individual journalists (BBC News, 2024). However, these responses have not yet resolved how GenAI should be integrated into collaborative newsroom practices: how teams of journalists and editors might coordinate GenAI use and maintain shared editorial standards in future collaborations.
@@
To understand why this missing layer of collaboration matters, we turn to the normative features of journalism that make it a particularly high-stakes and revealing site for studying GenAI’s integration into collaborative work. First, as a content-intensive profession, journalism is especially exposed to GenAI’s disruptions: core newsroom tasks such as brainstorming, reporting, and editing directly overlap with GenAI’s generative capabilities in text production, compelling journalists to negotiate how these systems intervene in their everyday collaborations (Cools and Diakopoulos, 2024). Second, newswork is inherently collaborative, structured through collective editorial routines and shared news values (Schudson, 2001; Deuze, 2005). GenAI therefore does not simply alter individual tasks but also potentially reshapes the collaborations and mutual trust on which newsroom collaboration depends. Third, the public visibility of journalism subjects GenAI adoption to heightened scrutiny, both inside and outside the newsroom. This visibility amplifies the stakes of collaborative decisions about when and how to use GenAI, since any misstep is not only organizationally consequential but also publicly accountable (Petre, 2021; Chadha and Koliska, 2016; Nishal and Diakopoulos, 2025; Møller, 2024; Komatsu et al., 2020).
@@
Yet despite journalism’s collaborative nature, much of the existing research on GenAI use has remained focused at the individual level. Early research on GenAI in newsrooms has largely focused on task-level affordances and risks for individuals. Scholars have highlighted how AI tools can accelerate content creation and support editorial work, while raising concerns about factual errors, bias amplification, content homogenization, and the erosion of core journalistic skills (Noh et al., 2025; Wang et al., 2025; Woodruff et al., 2024; Siitonen et al., 2024). More recent work has turned attention to value alignment and professional norms, showing that individual journalists adopt GenAI selectively and cautiously, favoring uses that preserve editorial credibility (Nishal and Diakopoulos, 2025; Tseng et al., 2025; Xiao et al., 2025b; Wu, 2024). However, a key gap remains: we know little about how GenAI becomes, or fails to become, a part of collaborative newsroom routines. By this we mean not sporadic or experimental uses of GenAI, but its sustained integration into everyday collaborative newswork, where its presence is taken for granted, embedded in daily teamwork, and aligned with collaborative norms.
@@
  • RQ1: What are journalists’ current experiences with GenAI that intersect with collaborative newsroom routines?
  • RQ2: What challenges do journalists face in integrating GenAI into their collaborative newsroom routines?
  • RQ3: How can newsrooms better integrate GenAI into collaborative newsroom routines in the future?
To answer these questions, we conducted an in-depth qualitative study of Chinese newsrooms. China is a fertile site for studying GenAI adoption in newsroom collaborations because Chinese newsrooms operate under both rapid technological transformation and strict professional expectations. Journalists navigate a dual pressure: leveraging cutting-edge GenAI tools to enhance innovation under governmental expectations, while upholding the editorial authority that underpins their professional norms (Xiao et al., 2025b). This dual pressure magnifies the challenges of integrating GenAI into collaborative newsroom routines and makes teamwork negotiation around GenAI especially visible.
Our study draws on 27 semi-structured interviews with senior newsroom managers, editors, and front-line journalists in China who had participated in early GenAI adoption in their newsrooms. Our findings showed that GenAI is already widely used at the individual level in Chinese newsrooms, but such use remains highly personalized, informal, and largely disconnected from collaborative routines (RQ1). Then, we found that structural fragmentation at the organizational level and cultural reluctance to share practices have prevented GenAI from being integrated into collaborative newsroom routines (RQ2). Journalists also envision GenAI-augmented collaboration as requiring a rethinking of organizational structures, cultural norms, and workflow integration (RQ3).
@@
  • First, we provide empirical evidence showing that GenAI adoption in newsrooms is currently individualized rather than collaborative. We highlight the gap between personal experimentation with GenAI usage and official organizational integration in collaborative work settings.
  • Second, we identify specific challenges that hinder GenAI integrations in newswork collaborations. We demonstrate how the absence of teamwork norms, coordination mechanisms, and shared practices impedes the emergence of accountable AI-augmented collaborative newsroom routines.
  • Third, we outline design and policy implications to inform the development of GenAI systems and organizational practices, fostering more effective integration of GenAI into collaborative work.
@@ 2. Related Work
We firstly review research on GenAI in the news industry (subsection 2.1). We then turn to broader studies of collaborations in the newsroom (subsection 2.2). Finally, we examine work on GenAI for teamwork and organizational collaboration (subsection 2.3).
@@ 2.1. GenAI in the News Industry
This subsection reviews prior work on GenAI in journalism, highlighting both its opportunities and risks (subsubsection 2.1.1) as well as the value tensions that potentially shape its cautious adoption in collaborative newsroom routines (subsubsection 2.1.2).
@@ 2.1.1. Opportunities and Concerns of GenAI in Journalism
Many news organizations now develop GenAI for drafting stories, summarizing reports, suggesting headlines, or even generating interview questions (Lewis et al., 2025; Ogola, 2024). Editing tasks are also supported through AI tools that propose alternative phrasings, check stylistic consistency, or flag potential bias (Noh et al., 2025; Wang et al., 2025). A 2023 global survey found that over 75% of news organizations across 46 countries were already using AI (JournalismAI, 2023). Some leading newsrooms such as China Media Group and Bloomberg are collaborating with AI technologists and experimenting with building tailored GenAI systems, seeking to customize GenAI for journalistic needs (Xiao et al., 2025b).
@@
Despite these affordances, concerns persist. Chief among these are the risks of factual inaccuracy, the amplification of existing biases, and the homogenization of content, where widespread reliance on the same tools may diminish the distinctive voice and editorial identity of individual outlets (Siitonen et al., 2024; Lewis et al., 2025; Thomson et al., 2024; Matich et al., 2025). Journalists have also voiced existential anxieties about de-skilling and professional displacement, fearing that GenAI may automate precisely those creative and investigative tasks that define the integrity and social value of their work  (Woodruff et al., 2024; Siitonen et al., 2024). As Woodruff et al. argue  (Woodruff et al., 2024), these threats are particularly acute given the journalism industry’s already-precarious position, marked by economic instability, declining public trust, and shrinking editorial capacity.
@@
HCI researchers have responded by exploring how GenAI tools can be aligned with newsroom practices (Aitamurto et al., 2019; Tseng et al., 2025; Xiao et al., 2025b; Nishal and Diakopoulos, 2025; Komatsu et al., 2020). For instance, Tseng et al.  (Tseng et al., 2025) proposed a journalist-centered design space for GenAI, arguing that systems must adapt to the editorial logics and workflows of the profession, rather than impose model-driven efficiencies. Hoque et al. (Hoque et al., 2024) designed a question-answering chatbot to assist journalists in presenting news content to readers in a balanced and objective manner.
2.1.2. Tensions Between GenAI and Newswork Routines
@@
Several studies have highlighted how journalistic collective values such as independence, transparency, fairness, and credibility function not just as abstract ideals, but as design requirements for GenAI systems in newsrooms and sources of sociotechnical friction (Nishal and Diakopoulos, 2025; Møller, 2024; Komatsu et al., 2020; Xiao et al., 2025b). Nishal and Diakopoulos (Nishal and Diakopoulos, 2025) argue that journalistic values must be treated as central constraints in system development. Other scholars emphasizes that editorial judgment is a situated professional norm that must be maintained even as automation expands in newswork routines (Xiao et al., 2025b; Møller, 2024). Many newsrooms also formalized their AI policies (Hofeditz et al., 2025; Viner and Bateson, [n. d.]). Major organizations such as The Guardian have published explicit editorial guidelines on GenAI, focusing on the core values of the newswork when using GenAI (Viner and Bateson, [n. d.]).
@@
These value tensions between GenAI and journalistic practice are not confined to top-down organizational directives. Recent studies show that journalists themselves are also experimenting with GenAI in informal, bottom-up ways (Wu, 2024). Wu (Wu, 2024) characterizes this as a “value-motivated use” framework, where journalists selectively adopt AI only insofar as it supports core editorial standards. In practice, individual journalists used GenAI for tasks such as designing interview questions, assisting with writing, clarifying terminology, summarizing key points, and checking facts. Yet these uses were marked by caution, as journalists had to constantly guard against potential errors, such as hallucinations, that could compromise their work (Wu, 2024). However, Wu (Wu, 2024) and other journalism studies do not examine how these individual practices relate to broader collaborative norms.
2.2. Collaborations in the Newsrooms
@@
One core form of collaboration is the front-line journalist–editor relationship. Front-line journalists generate ideas, gather information, and draft stories, while editors provide oversight, frame narratives, and ensure adherence to professional values such as accuracy, fairness, and credibility (Frost, 2010; Kolodzy, 2006; Duffy, 2021). Duffy (Duffy, 2021) argued that much of journalism research tends to conflate the roles of front-line journalists and editors, overlooking the crucial distinctions between them. While front-line journalists are often foregrounded as the visible producers of news, editors play a central role as the gatekeepers of news values.
Another dimension of collaboration occurs among front-line journalists themselves. Carlson (Carlson, 2009) points out that front-line journalists, even when affiliated with different news organizations, often end up working in shared spaces such as press conferences, media briefings, or public events. They sometimes also directly coordinate with one another in joint reporting arrangements (Dailey et al., 2005; Carlson, 2009). The collaborations of front-line journalists create conditions where information is rapidly exchanged and professional norms are reinforced through peer interactions (Crouse, 2003).
Collaboration in newsrooms increasingly spans both internal professional boundaries and external organizational ones. Internally, front-line journalists and editors work closely with photojournalists, data journalists, designers, and technical staff to produce multimedia stories, navigating tensions between journalistic norms and technical or esthetic priorities (Xiao et al., 2025b; Wu et al., 2021; Pavlik, 2019). For instance, Xiao et al. (Xiao et al., 2025b) documented how in-house AI engineers and journalists collaborated to build GenAI tools: engineers providing technical infrastructure, and journalists contributing normative and contextual expertise. Externally, some today’s journalists also need to negotiate collaborations with technology companies that supply GenAI systems (Dodds et al., 2025; Xiao et al., 2025c; Simon, 2024); such partnerships then raise new questions about editorial independence (Simon, 2024).

### v3

# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Coordinating GenAI into Collaborative Newswork

@@ Abstract.
Generative AI (GenAI) is reshaping work, but adoption remains largely individual and experimental rather than coordinated into collaborative work. Whether GenAI can move from individual use to collaborative work is a critical question for future organizations. Journalism offers a compelling site to examine this shift: individual journalists have already been disrupted by GenAI tools; yet newswork is inherently collaborative relying on shared norms and coordinated workflows. We conducted 27 interviews with newsroom managers, editors and front-line journalists in China. We found that journalists frequently used GenAI to support daily tasks, but value alignment was safeguarded mainly through individual discretion. At the organizational level, GenAI use remained disconnected from team workflows, hindered by structural barriers and cultural reluctance to share practices. These findings underscore the gap between individual and collaborative work, pointing to the need to account for organizational structures, cultural norms, and workflow when coordinating GenAI for collaborative work.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790984isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Existing discussions of GenAI adoption often assume that organizations play an active, top-down role in deciding when and how to integrate GenAI (Seifdar and Amiri, 2025; Xiao et al., 2025c). However, in many real-world settings, GenAI enters organizations not through coordinated institutional adoption but through dispersed, informal, and individualized practices (Waters-Lynch et al., 2025; Wu, 2024). Some individuals are already using GenAI inside organizations, often invisibly, creating pressures, ambiguities, and expectations for future collaborative coordination. Accordingly, our study asks how the widespread yet individualized GenAI use already occurring inside organizations creates new pressures for collaborative coordination and team workflow, and how organizations might respond by meaningfully coordinate GenAI into collaborative work under these conditions.
@@
In this paper, we use the newsroom as our primary empirical site to examine how individualized GenAI usage interacts with organizational collaboration. News production is rarely the effort of a single individual: front-line journalists gather information, draft stories, and exchange leads with peers; their work is then reviewed, edited, and sometimes rewritten by editors; and senior managers provide strategic oversight, shape policies, and liaise with external stakeholders (Duffy, 2021; Xiao et al., 2025c; Tuchman, 1973). Collaboration in newsrooms therefore spans both inter-role and intra-role interactions, anchored in shared norms and collective values such as editorial responsibility and mutual trust (Tuchman, 1973; Carlson, 2009; Robinson, 2011).
@@
While these roles structure collaboration in conventional newswork, the rise of GenAI introduces new uncertainties about how such collaboration could unfold. Recently, on the one hand, many individual journalists are experimenting with GenAI in tentative, ad hoc ways (Liang et al., 2025), often relying on personal judgment to ensure that its use remains consistent with professional norms (Wu, 2024). On the other hand, major news organizations have started to articulate institutional responses to the rise of GenAI, for example, China Media Group (CMG) is developing proprietary GenAI systems for its news production (China Press, Publication, Radio, Film and Television Journal, 2024; Lang, 2024), while The British Broadcasting Corporation (BBC) has issued formal usage guidelines for individual journalists (BBC News, 2024). However, these organizational responses do not yet address a crucial issue: how the already-prevalent individual uses of GenAI could be coordinated or aligned within collaborative newswork: how teams of journalists and editors might coordinate their GenAI use and maintain shared editorial standards in future collaborations.
@@
This gap between widespread individual use of GenAI and the absence of collaborative structures is the central problem this study addresses. To understand why this missing layer of collaboration matters, we turn to the normative features of journalism that make it a particularly high-stakes and revealing site for studying GenAI use in collaborative work. First, as a content-intensive profession, journalism is especially exposed to GenAI’s disruptions: core newsroom tasks such as brainstorming, reporting, and editing directly overlap with GenAI’s generative capabilities in text production, compelling journalists to negotiate how these systems intervene in their everyday collaborations (Cools and Diakopoulos, 2024). Second, newswork is inherently collaborative, structured through collective editorial routines and shared news values (Schudson, 2001; Deuze, 2005). GenAI therefore does not simply alter individual tasks but also potentially reshapes the collaborations and mutual trust on which newsroom collaboration depends. Third, the public visibility of journalism subjects GenAI adoption to heightened scrutiny, both inside and outside the newsroom. This visibility amplifies the stakes of collaborative decisions about when and how to use GenAI, since any misstep is not only organizationally consequential but also publicly accountable (Petre, 2021; Chadha and Koliska, 2016; Nishal and Diakopoulos, 2025; Møller, 2024; Komatsu et al., 2020).
@@
Yet despite journalism’s collaborative nature, much of the existing research on GenAI use has remained focused at the individual level. Early research on GenAI in newsrooms has largely focused on task-level affordances and risks for individuals. Scholars have highlighted how AI tools can accelerate content creation and support editorial work, while raising concerns about factual errors, bias amplification, content homogenization, and the erosion of core journalistic skills (Noh et al., 2025; Wang et al., 2025; Woodruff et al., 2024; Siitonen et al., 2024). More recent work has turned attention to value alignment and professional norms, showing that individual journalists adopt GenAI selectively and cautiously, favoring uses that preserve editorial credibility (Nishal and Diakopoulos, 2025; Tseng et al., 2025; Xiao et al., 2025c; Wu, 2024). However, a key gap remains: we know little about how GenAI becomes, or fails to become, coordinated into collaborative newsroom newswork. Here, we refer not to sporadic or experimental uses of GenAI, but to its potential for sustained integration into everyday collaborative newswork, where its presence is taken for granted, embedded in ongoing teamwork, and coordinated with shared professional norms.
@@
  • RQ1: How does the existing individual use of GenAI affect collaborative work in newsrooms?
  • RQ2: What challenges do journalists face in coordinating GenAI into their collaborative newswork?
  • RQ3: How do journalists envision future organizational mechanisms and system design that could better coordinate GenAI use in collaborative newswork?
To answer these questions, we conducted an in-depth qualitative study of Chinese newsrooms. China is a fertile site for studying GenAI’s impact in newsroom collaborations because Chinese newsrooms operate under both rapid technological transformation and strict professional expectations. Journalists navigate a dual pressure: leveraging cutting-edge GenAI tools to enhance innovation under governmental expectations, while upholding the editorial authority that underpins their professional norms (Xiao et al., 2025c). This dual pressure magnifies the challenges of coordinating GenAI into collaborative newswork and makes teamwork negotiation around GenAI especially visible.
Our study draws on 27 semi-structured interviews with senior newsroom managers, editors, and front-line journalists in China who had participated in early GenAI adoption in their newsrooms. Our findings showed that GenAI is already widely used at the individual level in Chinese newsrooms, but such use remains highly personalized, informal, and largely disconnected from collaborative newswork (RQ1). Then, we found that structural fragmentation at the organizational level and cultural reluctance to share practices have prevented GenAI from being coordinated into collaborative newswork. Current Chinese newsrooms have invested heavily in developing internal GenAI systems, yet have devoted far less attention to coordinating GenAI use within collaborative newswork (RQ2). Journalists also envision future coordination of GenAI into collaborative newswork as requiring a rethinking of organizational structures, cultural norms, and workflow (RQ3).
@@
  • First, we provide empirical evidence showing that GenAI use in newsrooms is currently individualized rather than collaborative. We highlight the gap between personal experimentation with GenAI usage and official organizational coordination in collaborative work settings.
  • Second, we identify specific challenges that hinder GenAI coordination in collaborative newswork. We demonstrate how the absence of teamwork norms, coordination mechanisms, and shared practices impedes the coordination of GenAI into collaborative newswork.
  • Third, we outline design and policy implications to inform the development of GenAI systems and organizational practices, fostering more effective coordination of GenAI into collaborative work.
@@ 2. Related Work
We firstly review research on GenAI in the news industry (subsection 2.1). We then turn to broader studies of collaborations in the newsroom (subsection 2.2). Then, we examine work on GenAI for teamwork and organizational collaboration (subsection 2.3). We conclude by discussing why individual GenAI use cannot be analytically isolated from organizational dynamics and introduce prior HCI and CSCW work on the organizational adoption of emerging technologies (subsection 2.4).
@@ 2.1. GenAI in the News Industry
This subsection reviews prior work on GenAI in journalism, highlighting both its opportunities and risks (subsubsection 2.1.1) as well as the value tensions that potentially shape its cautious adoption in collaborative newswork (subsubsection 2.1.2).
@@ 2.1.1. Opportunities and Concerns of GenAI in Journalism
Many news organizations now develop GenAI for drafting stories, summarizing reports, suggesting headlines, or even generating interview questions (Lewis et al., 2025; Ogola, 2024). Editing tasks are also supported through AI tools that propose alternative phrasings, check stylistic consistency, or flag potential bias (Noh et al., 2025; Wang et al., 2025). A 2023 global survey found that over 75% of news organizations across 46 countries were already using AI (JournalismAI, 2023). Some leading newsrooms such as China Media Group and Bloomberg are collaborating with AI technologists and experimenting with building tailored GenAI systems, seeking to customize GenAI for journalistic needs (Xiao et al., 2025c).
@@
Despite these affordances—functional possibilities that GenAI makes available for journalistic work—concerns persist. Chief among these are the risks of factual inaccuracy, the amplification of existing biases, and the homogenization of content, where widespread reliance on the same tools may diminish the distinctive voice and editorial identity of individual outlets (Siitonen et al., 2024; Lewis et al., 2025; Thomson et al., 2024; Matich et al., 2025). Journalists have also voiced existential anxieties about de-skilling and professional displacement, fearing that GenAI may automate precisely those creative and investigative tasks that define the integrity and social value of their work  (Woodruff et al., 2024; Siitonen et al., 2024).
@@
HCI researchers have responded by exploring how GenAI tools can be aligned with newsroom practices (Aitamurto et al., 2019; Tseng et al., 2025; Xiao et al., 2025c; Nishal and Diakopoulos, 2025; Komatsu et al., 2020). For instance, Tseng et al.  (Tseng et al., 2025) proposed a journalist-centered design space for GenAI, arguing that systems could adapt to the editorial logics and workflows of the profession, rather than impose model-driven efficiencies. Hoque et al. (Hoque et al., 2024) designed a question-answering chatbot to assist journalists in presenting news content to readers in a balanced and objective manner.
2.1.2. Tensions Between GenAI and Newswork
@@
Several studies have highlighted how journalistic collective values such as independence, transparency, fairness, and credibility function not just as abstract ideals, but as design requirements for GenAI systems in newsrooms and sources of sociotechnical friction (Nishal and Diakopoulos, 2025; Møller, 2024; Komatsu et al., 2020; Xiao et al., 2025c). Nishal and Diakopoulos (Nishal and Diakopoulos, 2025) argue that journalistic values could be treated as central constraints in system development. Other scholars emphasizes that editorial judgment is a situated professional norm that could be maintained even as automation expands in newswork routines (Xiao et al., 2025c; Møller, 2024). Many newsrooms also formalized their AI policies (Hofeditz et al., 2025; Viner and Bateson, 2023). Major organizations such as The Guardian have published explicit editorial guidelines on GenAI, focusing on the core values of the newswork when using GenAI (Viner and Bateson, 2023).
@@
Recent studies show that journalists themselves are also experimenting with GenAI in informal, bottom-up ways (Wu, 2024). Wu (Wu, 2024) characterizes this as a “value-motivated use” framework, where journalists selectively adopt AI only insofar as it supports core editorial standards. In practice, individual journalists used GenAI for tasks such as designing interview questions, assisting with writing, clarifying terminology, summarizing key points, and checking facts. Yet these uses were marked by caution, as journalists had to constantly guard against potential errors, such as hallucinations, that could compromise their work (Wu, 2024). However, Wu (Wu, 2024) and other journalism studies do not examine how these individual practices relate to broader collaborative norms.
2.2. Collaboration in the Newsrooms
@@
One core form of collaboration is the front-line journalist–editor relationship. Front-line journalists generate ideas, gather information, and draft stories, while editors provide oversight, frame narratives, and ensure adherence to professional values such as accuracy, fairness, and credibility (Frost, 2010; Kolodzy, 2006; Duffy, 2021). Another dimension of collaboration occurs among front-line journalists themselves. Carlson (Carlson, 2009) points out that front-line journalists, even when affiliated with different news organizations, often end up working in shared spaces such as press conferences, media briefings, or public events (Dailey et al., 2005; Carlson, 2009). The collaborations of front-line journalists create conditions where information is rapidly exchanged and professional norms are reinforced through peer interactions (Crouse, 2003).
Collaboration in newsrooms increasingly spans both internal professional boundaries and external organizational ones. Internally, front-line journalists and editors work closely with photojournalists, data journalists, designers, and technical staff to produce multimedia stories, navigating tensions between journalistic norms and technical or esthetic priorities (Xiao et al., 2025c; Wu et al., 2021; Pavlik, 2019). For instance, Xiao et al. (Xiao et al., 2025c) documented how in-house AI engineers and journalists collaborated to build GenAI tools: engineers providing technical infrastructure, and journalists contributing normative and contextual expertise. Externally, some today’s journalists also need to negotiate collaborations with technology companies that supply GenAI systems (Dodds et al., 2025; Xiao et al., 2025d; Simon, 2024); such partnerships then raise new questions about editorial independence (Simon, 2024).